a boats crew rowed 18 miles downstream with the current in 2 hrs the return trip upstream against the current covered the same distance but took 4.5 hours. find the crews rowing rate in still water and the rate of the current.
the crew rowing rate in still water is --------- m.p.h
the rate of the current is -------------m.p.h

Answer by mananth(16949)

a boats crew rowed 18 miles downstream with the current in 2 hrs the return trip upstream against the current covered the same distance but took 4.5 hours. find the crews rowing rate in still water and the rate of the current.
the crew rowing rate in still water is --------- m.p.h
the rate of the current is -------------m.p.h
...
let speed of boat be x mph
let current speed be y mph
..
upstream
distance = 18 miles
speed = x-y mph
time = 18/(x-y)= 4.5hours
18=4.5(x-y)
divide by 4.5
x-y =4 ..............................1
...
downstream
distance = 18 miles
speed = x+y
time = 18/(x+y)=2
18=2(x+y)
/2
x+y=9................................2
add 1 & 2
2x=13
/2
x=6.5 mph ..... speed of boat
plug value of x in equation 1
we get y= 2.5 mph..... speed of current
...
m.ananth@hotmail.ca